A stack of 100 nickels is 6.25 inches high. To the nearest cent, how much would a stack of nickels 8 feet high be worth?
Answer by checkley71(8403)

100*.25=$25 the value of 100 nickels that are 6.25 inches high.
6.25/(8*12)=25/x
6.25x=25*96
6.25x=2400
x=2400/6.25
x=$384 worth of nickels.